the length of of a triangle are x cm ,{x+1} and {x+2}cm determine x so that this triangle is the right angled triangle .
thank you in advance
this question is from the book mathematics 6th edition written by teh keng seng

Answer by edjones(8007)

a^2+b^2=c^2
x^2+(x+1)^2=(x+2)^2
x^2+x^2+2x+1=x^2+4x+4
x^2-2x-3=0
(x-3)(x+1)=0
x=3
.
Ed